Use the normal distribution of SAT critical reading scores for which the mean is

511511

and the standard deviation is

106106.

Assume the variable x is normally distributed.

left parenthesis a right parenthesis(a)

What percent of the SAT verbal scores are less than

600600​?

left parenthesis b right parenthesis(b)

If 1000 SAT verbal scores are randomly​ selected, about how many would you expect to be greater than

525525​?
Answer #1

\mu = 511

\sigma = 106

(a)
To find P(X<600):

Z = (600 - 511)/106 = 0.8396

Table of Area Under Standard Normal Curve gives area = 0.2995

So,

P(X<600) = 0.5 + 0.2995 = 0.7995

(b)

To find P(X>525):
Z = (525 - 511)/106 = 0.1321

Table gives area = 0.0517

So,

P(X>525) = 0.5 -0.0517 = 0.4483

For 1000 SAT verbal scores, Number greater than 525 = 0.4483 X 1000 = 448

So,

Answers to questions asked:

(a) 0.7995
(b) 448